OpenStudy (anonymous):

describe the relationship between the graphs of the linear equation y=1/3x+10 and 3y=9x+2

OpenStudy (anonymous):

did you type the question wrong? it looks like it's going for a perpendicular relationship, but that would be a negative reciprocal if not, the second line is 9 times as steep (slope wise) as the first... and they would intercept at 1/3x + 10 = 3x + 2/3 x + 30 = 9x + 2 -8x = -28 x = 7/2 y = 335 / 30 = 67 / 6 intersect at (7/2, 67/6)

OpenStudy (anonymous):

the equations are: y=-1/3x +10 and 3y=9x-2

OpenStudy (anonymous):

i forgot a (-)

OpenStudy (anonymous):

it sounds like you know what you are doing though! lol

OpenStudy (anonymous):

lols then i was right in the first sentence. ignore the rest it's perpendicular

OpenStudy (anonymous):

i am just not sure how to solve the equation...how do i start?

OpenStudy (anonymous):

the definition of perpendicular lines requires that the slopes of the lines have to be negative reciprocals. Negative reciprocals means that it's negative and inversed (1/x) so in your example y=1/3x+10 and 3y=9x+2 change both to y = mb + c y=-1/3x+10 y=3x+2/3 so the slopes are -1/3 and 3 if you take the negative reciprocal of -1/3 would be \[\frac{1}{-\frac{1}{3}}* -1\] which would equal 3 Since they are negative reciprocals, they are perpendicular it's a rule memorize it just like how parallel lines have the same slope, it's the same idea perpendicular lines are negative reciprocals cheers ^_^

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no prob, good lucks